OpenStudy (anonymous):

The radius of a water molecule is about 1.4 angstroms. One angstrom is 0.00000001 cm. What is the diameter of a water molecule in centimeters? Use scientific notation! Please show your work!! Thanks!

OpenStudy (anonymous):

@MissNoName

OpenStudy (trojanpoem):

Diameter :D , Raduis : R D = 2R R = 1.4 angstrom 1 angstrom = 0.00000001 cm = 10^-7 cm R = 1.4 * 10^-7 cm D = 2.8 * 10^-7 cm

OpenStudy (anonymous):

Thanks!